# TopoTune: A Framework for Generalized Combinatorial Complex Neural Networks

### v3

@@ 7 Conclusion
This work introduces a simple yet powerful graph-based methodology for constructing Generalized Combinatorial Complex Neural Networks (GCCNs), TDL architectures that generalize and subsume standard CCNNs. Additionally, we introduce TopoTune, the first lightweight software module for systematically and easily implementing new TDL architectures across topological domains. In doing so, we have addressed, either in part or in full, 7 of the 11 open problems of the field defined by some of its leaders in Papamarkou et al. (2024). Future work includes customizing GCCNs for application-specific and potentially sparse or multimodal datasets, and leveraging software from state-of-the-art GNNs. TopoTune will also help bridge the gap with other fields such as attentional learning and $k$-hop higher-order GNNs (Morris et al., 2019; Maron et al., 2019).
@@ B.3 Proof of Expressivity
We provide the theory required to prove Proposition 4.3, i.e., to prove that GCNNs are strictly more expressive than CCNNs. The definitions and propositions from this subsection are summarized in Figure 7. This figure serves as a graphical reading guide for the subsection.
Figure 7: Graphical summary of the definitions and propositions related to the expressivity of CCNNs and GCNNs and of the different WL tests. Neural networks expressivity is in red, and WL test expressivity is in blue.
@@ B.3.1 Homomorphism and Isomorphism Induced by Neighborhoods
We first recall the notion of homomorphism of a combinatorial complexes (CC) from (Hajij et al., 2023) and generalize it to the notions of homomorphism and of isomorphism of combinatorial complexes induced by a neighborhood $\mathcal{N}$.
@@ B.3.3 Definitions of $k$-GNNs and $k$-CCNNs
Then, we show that $k$-CCNNs of Definition B.9 form a subclass of GCNNs.
@@
GCNNs generalize and subsume $k$-CCNNs.
@@
Let be given a $k$-CCNN defined by (11). We show that we can recover (11) by an appropriate choice of feature dimensionality $F$, update function $\phi$ and sub-module $\omega_{\mathcal{N}}$ in (8) defining GCNNs, and thus that any $k$-CCNN can be expressed as a GCNN.
For simplicity of notations, we assume that the layers of the $k$-CCNN have same feature dimensionality, denoted $d$. Given that there are $\binom{|\mathcal{C}|-1}{k-1}$ $k$-sets containing a given cell $\sigma$, we define $F=\binom{|\mathcal{C}|-1}{k-1}d$ to be the feature dimensionality of the layers of a GCNN. Denote $\mathbf{H}^{l}(\sigma)$ the row of $\mathbf{H}^{l}$ containing $F$-dimensional feature corresponding to cell $\sigma$, as well as $\mathbf{H}^{l}(\sigma,s)$ the subrow containing the $d$-dimensional feature corresponding to one $k$-set $s$ to which $\sigma$ belongs. Figure 8 illustrates these notations. We then define:
@@
In other words, we first use $\omega$ defined as a sequence of $\omega_{(\sigma,s)}$ to update each $(\sigma,s)$-block of $\mathbf{H}^{l}$ into an auxiliary feature $B=\tilde{\mathbf{H}}^{l}$. Then, we use $\phi$ as a sequence of $\phi_{(\sigma,s)}$ to perform a block-wise operations. Thus, we have built a GCNN that reproduces the computations of the $k$-CCNN. Therefore, GCNNs generalize and subsume $k$-CCNNs. ∎
B.3.4 Relationships between CCWL/GCWL tests and CCNNs/GCNNs
@@
Theorem 1 in (Morris et al., 2019) states that, for every encoding of the graph labels $\ell(v)$ as $d$-vectors $h^{(0)}(v)$, and for every choice of parameters $W^{(t)}$, the coloring $c(t)_{\ell}$ of the WL test always refines the coloring $h(t)$ induced by the GNN parameterized by $W^{(t)}$. Theorem 2 in (Morris et al., 2019) states that there exists parameter matrices $W^{(t)}$ such that GNNs have exactly the same power as the WL test. Consequently, we say that GNNs have the same expressivity as the WL test. Similarly, Propositions 3 and 4 from (Morris et al., 2019) show that $k$-GNNs have the same expressivity as the WL test.
CCWL/CCNNs and $k$-CCWL/GCNNs
@@
Consequently, CCNNs have the same power as the CCWL. Next, by Proposition B.12, we can also pull the results from Propositions 3 and 4 from (Morris et al., 2019) and provide the following propositions.
@@
Let $(\mathcal{C},\ell)$ be a labeled CC and let $k\geq 2$. Then, for all $t\geq 0$, for all choices of initial colorings $h_{k}^{(0)}$ consistent with $\ell$ and for all weights $\mathbf{W}^{(t)}$, $c_{\mathrm{s},k,\ell}^{(t)}\sqsubseteq h_{k}^{(t)}.$
@@
Consequently, $k$-CCNNs have the same power as the $k$-CCWL.
@@ B.3.5 Proof
We prove that GCNNs are strictly more powerful than CCNNs in distinguishing non-isomorphic combinatorial complexes. We leverage the propositions of this subsection summarized on Figure 7.
By Proposition B.10, GCNN have at least the same expressive power as $k$-CCNNs. By Propositions B.15-B.16, $k$-CCNNs have the same expressive power as the $k$-CCWL. By Proposition B.12, the $k$-CCWL test is equivalent to the $k$-WL test on the associated strictly augmented Hasse graph. It is known (e.g., (Grohe, 2017)) that the $k$-WL test on graph is strictly more powerful than the WL test. Thus, the $k$-WL test on the strictly augmented Hasse graph is strictly more powerful than the WL test on that same graph. By Proposition B.11, the WL test on the strictly augmented Hasse graph is equivalent to the CCWL test on the corresponding CC. By Propositions B.13-B.14, the CCWL test on the CC has the same expressive power as CCNNs.
Consequently, we have shown that GCNN are strictly more powerful than CCNNs in distinguishing nonisomorphic CCs. ∎
@@
Additionally, we construct two combinatorial complexes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ that are indistinguishable by CCNNs but distinguishable by GCNNs.
@@
Since $k$-CCWL is equivalent to $k$-WL on $\mathcal{G}_{\mathcal{N}}$ (Proposition B.12), the two complexes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are distinguishable by $k$-CCWL. Since $k$-CCWL has the same expressive power as $k$-CCNNs (Propositions B.15-B.16), the two complexes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are distinguishable by $k$-CCNNs. Since GCNNs generalize and subsume $k$-CCNNs (Proposition B.10), $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are distinguishable by GCNNs.
Thus, we have constructed two combinatorial complexes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ that are indistinguishable by CCNNs, but are distinguishable by GCNNs.
@@ Appendix D Software
Remark. We decided to design the software module of TopoTune, i.e., how to implement GCCNs, as we did for mainly two reasons: (i) the full compatibility with TopoBenchmark (implying consistency of the combinatorial complex instantiations and the benchmarking pipeline), and (ii) the possibility of using GNNs as neighborhood message functions that are not necessarily implemented with a specific library. However, if the practitioner is interested in entirely wrapping the GCCN implementation into Pytorch Geometric or DGL, they can do it by noticing that a GCCN is equivalent to a heterogeneous GNN where the heterogeneous graph the whole augmented Hasse graph, with node types given by the rank of the cell (e.g. 0-cells, 1-cells, and 2-cells) while the edge type is given by the per-rank neighborhood function (e.g. ”0-cells to 1-cells” or ”2-cells to 1-cells” for $\mathcal{N}^{0}_{I,\uparrow}$ and $\mathcal{N}^{2}_{I,\downarrow}$, respectively).
@@ E.4 Hardware
Appendix F Iterating upon existing CCNNs
We design and test GCCNs that replicate the neighborhood structure and training scheme of existing CCNNs, and replace their message-passing schemes with the same GNNs considered in Table 1. Results in Table 3 show that this sometimes leads to better performance on benchmark datasets.
Model MUTAG PROTEINS NCI1 NCI109 Cora Citeseer PubMed
SCCN (Yang et al., 2022)
CWN (Bodnar et al., 2021a)
Appendix G Model Size
@@
CCNN (Best Model on TopoBenchmark) 334.72K 101.12K 63.87K 17.67K 88.06K 451.85K 1032.84K 163.72K
@@
CCNN (Best Model on TopoBenchmark) 398.85K 10.24K 131.84K 135.75K 617.86K 144.62K 737.29K 134.40K
@@
CCNN (Best Model on TopoBenchmark) 84.10K 14.34K 88.19K 88.32K 22.53K 60.26K 258.50K 280.83K
@@
Appendix H Model Training Time
@@
CCNN (Best Model on TopoBenchmark) 100 ± 23 132 ± 19 238 ± 89 254 ± 39 228 ± 44 75 ± 15 57 ± 4.4 128 ± 50
@@
CCNN (Best Model on TopoBenchmark) 123 ± 57 104 ± 28 172 ± 50 183 ± 62 178 ± 86 143 ± 16 75 ± 23 114 ± 18
@@
CCNN (Best Model on TopoBenchmark) 127 ± 48 96 ± 20 220 ± 74 128 ± 49 387 ± 105 121 ± 38 48 ± 1 177 ± 71

### v4

@@ 7 Conclusion
This work introduces a simple yet powerful graph-based methodology for constructing Generalized Combinatorial Complex Neural Networks (GCCNs), TDL architectures that generalize and subsume standard CCNNs. Additionally, we introduce TopoTune, the first lightweight software module for systematically and easily implementing new TDL architectures across topological domains. In doing so, we have addressed, either in part or in full, 7 of the 11 open problems of the field (Papamarkou et al., 2024). Future work includes customizing GCCNs for application-specific and potentially sparse or multimodal datasets, and leveraging software from state-of-the-art GNNs. TopoTune will also help bridge the gap with other fields such as attentional learning and $k$-hop higher-order GNNs (Morris et al., 2019; Maron et al., 2019).
Acknowledgements
M.P. acknowledges the support of National Science Foundation (NSF) CAREER 2240158 and NSF Grant 2134241, as well as from the National Science and Engineering Research Council of Canada. G.B. acknowledges support from NSF Grant 2134241. C.B. acknowledges support from the National Institutes of Health Grant 1R01ES037156-01. N.M. acknowledges support from NSF Grant 2313150.
@@ B.3 Proof of Expressivity
We provide the theory required to prove Proposition 4.3, i.e., to prove that GCCNs are strictly more expressive than CCNNs. The definitions and propositions from this subsection are summarized in Figure 7. This figure serves as a graphical reading guide for the subsection.
Figure 7: Graphical summary of the definitions and propositions related to the expressivity of CCNNs and GCCNs and of the different WL tests. Neural networks expressivity is in red, and WL test expressivity is in blue.
@@ B.3.1 Homomorphism and Isomorphism Induced by Neighborhoods
We first recall the notion of homomorphism of a combinatorial complex (CC) from (Hajij et al., 2023) and generalize it to the notions of homomorphism and isomorphism of CCs induced by a neighborhood $\mathcal{N}$.
@@ B.3.3 Definitions of $k$-GNNs and $k$-CCNNs
Then, we show that $k$-CCNNs of Definition B.9 form a subclass of GCCNs.
@@
GCCNs generalize and subsume $k$-CCNNs.
@@
Let be given a $k$-CCNN defined by (11). We show that we can recover (11) by an appropriate choice of feature dimensionality $F$, update function $\phi$ and sub-module $\omega_{\mathcal{N}}$ in (8) defining GCCNs, and thus that any $k$-CCNN can be expressed as a GCCN.
For simplicity of notations, we assume that the layers of the $k$-CCNN have same feature dimensionality, denoted $d$. Given that there are $\binom{|\mathcal{C}|-1}{k-1}$ $k$-sets containing a given cell $\sigma$, we define $F=\binom{|\mathcal{C}|-1}{k-1}d$ to be the feature dimensionality of the layers of a GCCN. Denote $\mathbf{H}^{l}(\sigma)$ the row of $\mathbf{H}^{l}$ containing $F$-dimensional feature corresponding to cell $\sigma$, as well as $\mathbf{H}^{l}(\sigma,s)$ the subrow containing the $d$-dimensional feature corresponding to one $k$-set $s$ to which $\sigma$ belongs. Figure 8 illustrates these notations. We then define:
@@
In other words, we first use $\omega$ defined as a sequence of $\omega_{(\sigma,s)}$ to update each $(\sigma,s)$-block of $\mathbf{H}^{l}$ into an auxiliary feature $B=\tilde{\mathbf{H}}^{l}$. Then, we use $\phi$ as a sequence of $\phi_{(\sigma,s)}$ to perform a block-wise operations. Thus, we have built a GCCN that reproduces the computations of the $k$-CCNN. Therefore, GCCNs generalize and subsume $k$-CCNNs. ∎
B.3.4 Relationships between CCWL/GCWL tests and CCNNs/GCCNs
@@
Theorem 1 in (Morris et al., 2019) states that, for every encoding of the graph labels $\ell(v)$ as $d$-vectors $h^{(0)}(v)$, and for every choice of parameters $W^{(t)}$, the coloring $c(t)_{\ell}$ of the WL test always refines the coloring $h(t)$ induced by the GNN parameterized by $W^{(t)}$. Theorem 2 in (Morris et al., 2019) states that there exists parameter matrices $W^{(t)}$ such that GNNs have exactly the same power as the WL test. Consequently, we say that GNNs have the same expressivity as the WL test. Similarly, Propositions 3 and 4 from (Morris et al., 2019) show that $k$-GNNs have the same expressivity as the k-WL test.
CCWL/CCNNs and $k$-CCWL/GCCNs
@@
Consequently, CCNNs have the same power as the CCWL. Next, we measure the power of $k$-CCNNs using the $k$-CCWL.
@@
Let $(\mathcal{C},\ell)$ be a labeled CC and let $k\geq 2$. Then, for all $t\geq 0$, for all choices of initial colorings $h_{k}^{(0)}$ consistent with $\ell$ and for all weights $\mathbf{W}^{(t)}$, $c_{\mathrm{s},k,\ell}^{(t)}\sqsubseteq h_{k}^{(t)}$ i.e., the coloring $c_{s,k,l}^{(t)}$ induced by the $k$-CCWL test refines the coloring induced by the $k$-CCNN $h_{k}^{(t)}$. .
@@
Propositions B.15 and B.16 are given by Proposition B.12 and Propositions 3 and 4 from Morris et al. (2019). They show that $k$-CCNNs have the same power as the $k$-CCWL.
@@ B.3.5 Proof
We prove that GCCNs are strictly more powerful than CCNNs in distinguishing non-isomorphic combinatorial complexes. We leverage the propositions of this subsection summarized on Figure 7.
By Proposition B.10, GCCN have at least the same expressive power as $k$-CCNNs. By Propositions B.15-B.16, $k$-CCNNs have the same expressive power as the $k$-CCWL. By Proposition B.12, the $k$-CCWL test is equivalent to the $k$-WL test on the associated strictly augmented Hasse graph. It is known (e.g., (Grohe, 2017)) that the $k$-WL test on graph is strictly more powerful than the WL test. Thus, the $k$-WL test on the strictly augmented Hasse graph is strictly more powerful than the WL test on that same graph. By Proposition B.11, the WL test on the strictly augmented Hasse graph is equivalent to the CCWL test on the corresponding CC. By Propositions B.13-B.14, the CCWL test on the CC has the same expressive power as CCNNs.
Consequently, we have shown that GCCN are strictly more powerful than CCNNs in distinguishing nonisomorphic CCs. ∎
@@
Additionally, we construct two combinatorial complexes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ that are indistinguishable by CCNNs but distinguishable by GCCNs.
@@
Since $k$-CCWL is equivalent to $k$-WL on $\mathcal{G}_{\mathcal{N}}$ (Proposition B.12), the two complexes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are distinguishable by $k$-CCWL. Since $k$-CCWL has the same expressive power as $k$-CCNNs (Propositions B.15-B.16), the two complexes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are distinguishable by $k$-CCNNs. Since GCCNs generalize and subsume $k$-CCNNs (Proposition B.10), $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are distinguishable by GCCNs.
Thus, we have constructed two combinatorial complexes $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ that are indistinguishable by CCNNs, but are distinguishable by GCCNs.
@@ Appendix D Software
Remark. We decided to design the software module of TopoTune, i.e., how to implement GCCNs, as we did for mainly two reasons: (i) the full compatibility with TopoBench (implying consistency of the combinatorial complex instantiations and the benchmarking pipeline), and (ii) the possibility of using GNNs as neighborhood message functions that are not necessarily implemented with a specific library. However, if the practitioner is interested in entirely wrapping the GCCN implementation into Pytorch Geometric or DGL, they can do it by noticing that a GCCN is equivalent to a heterogeneous GNN where the heterogeneous graph the whole augmented Hasse graph, with node types given by the rank of the cell (e.g. 0-cells, 1-cells, and 2-cells) while the edge type is given by the per-rank neighborhood function (e.g. ”0-cells to 1-cells” or ”2-cells to 1-cells” for $\mathcal{N}^{0}_{I,\uparrow}$ and $\mathcal{N}^{2}_{I,\downarrow}$, respectively).
@@ E.4 Hardware
Appendix F Model Size
@@
CCNN (Best Model on TopoBench) 334.72K 101.12K 63.87K 17.67K 88.06K 451.85K 1032.84K 163.72K
@@
CCNN (Best Model on TopoBench) 398.85K 10.24K 131.84K 135.75K 617.86K 144.62K 737.29K 134.40K
@@
CCNN (Best Model on TopoBench) 84.10K 14.34K 88.19K 88.32K 22.53K 60.26K 258.50K 280.83K
@@
Graph-Level Tasks Node-Level Tasks
@@
Appendix G Model Training Time
@@
CCNN (Best Model on TopoBench) 100 ± 23 132 ± 19 238 ± 89 254 ± 39 228 ± 44 75 ± 15 57 ± 4.4 128 ± 50
@@
CCNN (Best Model on TopoBench) 123 ± 57 104 ± 28 172 ± 50 183 ± 62 178 ± 86 143 ± 16 75 ± 23 114 ± 18
@@
CCNN (Best Model on TopoBench) 127 ± 48 96 ± 20 220 ± 74 128 ± 49 387 ± 105 121 ± 38 48 ± 1 177 ± 71
@@
Graph-Level Tasks Node-Level Tasks
@@
Appendix H Additional experiments
H.1 Larger node-level datasets
Table 8 additionally presents the experimental results on 4 heterophilic datasets introduced in (Platonov et al., ) (Amazon Ratings, Roman Empire, Minesweeper, and Questions). These represent larger node-level classification tasks than those shown in the main Table 1, with up to 48,921 nodes and 153,540 edges in the case of the Questions graph. Except on this precise dataset, which was not considered in previous TDL literature, we compare the results against CCNNs and hypergraph models from Telyatnikov et al. (2024). We observe that overall GCCNs achieve similar performance than regular CCNNs, and they outperform them by a significant margin on Minesweeper.
Amazon Ratings Roman Empire Minesweeper Questions
Best GCCN Cell 50.17 ± 0.71 84.48 ± 0.29 94.02 ± 0.28 78.04 ± 1.34
Best CCNN Cell 51.90 ± 0.15 82.14 ± 0.00 89.42 ± 0.00 -
Best GCCN Simplicial 50.53 ± 0.64 88.24 ± 0.51 94.06 ± 0.32 77.43 ± 1.33
Best CCNN Simplicial OOM 89.15 ± 0.32 90.32 ± 0.11 -
Best Hypergraph Model 50.50 ± 0.27 81.01 ± 0.24 84.52 ± 0.05 -
Table 8: Results on larger node level datasets, each experiment run with 5 seeds. We report accuracy for Amazon Ratings and Roman Empire, and AUC-ROC for Minesweeper and Questions. The values for the best CCNNs and hypergraph models are extracted from TopoBench (Telyatnikov et al., 2024).
H.2 More advanced GNNs
We include a subset of experiments with the same protocol as in Table 1 using GATv2 (Brody et al., 2021) and PNA (Corso et al., 2020) in the cellular domain. Results show how the GCCNs built with these models perform consistently well across node-level and graph-level tasks on the cell domain, often $<1\sigma$ of the best (standard-GNN) GCCN as reported in Table 1, but only outperform them on MUTAG.
Table 9: Cross-domain, cross-task, cross-expansion, and cross-$\omega_{\mathcal{N}}$ comparison of GCCN architectures built with GATv2 (Brody et al., 2021) and PNA (Corso et al., 2020) with top-performing GCCNs from Table 1 and benchmarked on TopoBench (Telyatnikov et al., 2024). Best result is in bold and results within 1 standard deviation are highlighted blue. Experiments are run with 5 seeds. We report accuracy for classification tasks and MAE for regression.
Graph-Level Tasks Node-Level Tasks
Model MUTAG ($\uparrow$) PROTEINS ($\uparrow$) NCI1 ($\uparrow$) NCI109 ($\uparrow$) Cora ($\uparrow$) Citeseer ($\uparrow$) PubMed ($\uparrow$)
Cellular
CCNN (Best Model on TopoBench) 80.43 ± 1.78 76.13 ± 2.70 76.67 ± 1.48 75.35 ± 1.50 87.44 ± 1.28 75.63 ± 1.58 88.64 ± 0.36
GCCN $\omega_{\mathcal{N}}$ = Best Standard GNN 86.38 ± 6.49 74.41 ± 1.77 78.23 ± 1.47 77.10 ± 0.83 88.57 ± 0.58 75.89 ± 1.84 89.40 ± 0.57
GCCN $\omega_{\mathcal{N}}$ = PNA 83.83 ± 6.31 73.91 ± 2.63 77.24 ± 1.72 76.5 ± 0.88 87.27 ± 0.64 74.63 ± 1.32 86.34 ± 0.38
GCCN $\omega_{\mathcal{N}}$ = GATv2 86.38 ± 4.15 72.54 ± 3.3 77.78 ± 0.94 77.04 ± 0.63 85.11 ± 0.46 72.03 ± 2.54 88.32 ± 0.38